OpenStudy (anonymous):

Which statement describes the translation of the graph of y = 3(x + 1)^2 – 1 from standard position? A. Moved down and to the right. B. Moved up and to the right. C. Moved down and to the left. D. Moved up and to the left.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

you see why its down ?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Moved down and to the left?

OpenStudy (anonymous):

lets take for example : y = x^2 - 1 incompare to y = x^2 when you plug x=0 into y =x^2 you get y =0 but if you plug x =0 into y = x^2 - 1 you get y = -1

OpenStudy (anonymous):

how'd u figure it out for this 1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

treat the whole term 3(x + 1)^2 as some function x^2 for example so its the same

OpenStudy (anonymous):

it is always C

OpenStudy (anonymous):

:O